KINETICS OF VIBRATIONAL UP-PUMPING STUDIED BY STEP-SCAN FOURIER TRANSFORM SPEC-TROSCOPY

\underline{E. PLÖNJES ^{a}}, P. PALM ^{a}, I. ADAMOVICH, J.W. RICH, Department of Mechanical Engineering, The Ohio State University, Columbus, OH 43210; W. URBAN, Institut für Angewandte Physik der Universität Bonn, 53115 Bonn, Germany.

Gaseous CO in an Ar diluent is prepared in vibrational state v = 1 by absorption of pulsed CO laser radiation. Energy subsequently diffuses into higher vibrational states ( to v = 40) by vibration-to-vibration exchange pumping. This diffusion of energy through vibrational quantum states is observed by step-scan FTIR emission spectroscopy on the 1^{st} overtone CO bands. The method permits monitoring of the time evolution of the vibrational state populations during the exchange pumping process. Δv = 2 emission spectra at 1 cm^{−1} resolution and 50 μsec time intervals are obtained. Vibrational stateresolved energy exchange rates are inferred from these measurements. In addition, state-resolved associative ionization rates for collisions of vibrationally excited CO are inferred.

aVisiting Scholar at The Ohio State University, permanently at the Institut für Angewandte Physik der Universität Bonn, 53115 Bonn, Germany